Dublin University Rifle Club

Rhona Barry, Ciara Walsh, Janet Chua and Derval Tubridy

Do you want to try a sport that caters for all ages, ablilities and ambitions?

Target shooting is a sport where brawn is not required to achieve success. Men and women can compete equally without handicap. You can compete against clubs from all over the world without having to leave the campus and you can enjoy a healthy social community at all levels from the purely recreational to competing in the olympics. Interested ?

Then welcome to the wonderful world of target shooting !

Contacting the Club

We have a noticeboard in Front Arch, marked on the map. All competition notices are published there at least a week in advance. We run an E-mail mailing list, to join simply send an E-mail to Majordomo@tcd.ie with subscribe durc-list in the message. All information posted on the noticeboard is sent out on the mailing list at the same time. The club has it’s own website at http://durc.tcd.ie/ where you can find notices of events, results, general news and photographs of various events. We’re planning on adding to this website during the year to add things like bulletin-board discussion forums, webcam broadcasts of competitions and on-line booking of shooting details. Our phone number is +353 − 1 − 608 − 1281. You can send regular mail to :

Dublin University Rifle Club c/o DUCAC, Trinity College Dublin, Dublin 2.

Joining the Club

So you’re interested ? Well you’re welcome to join us at any time. It costs $3 to join the club for the year. Come down to the range during term and sign up, we’re open from 1800 onwards, Monday through Friday. But note that you have to join up before you can book a detail to shoot!

Safety Meeting

We hold a meeting for Safety and Orientation every year. We go though the safety rules and precautions, demonstrate the rifles and shooting positions, and explain how to book details and so forth. You can also collect your membership cards at the meeting, and afterwards there are drinks and snacks in the clubhouse.

WHEN IS THE MEETING ON ???

This year the meeting is in

Davis Theatre, Arts Building

at

1800 on Monday, October 8th

About Target Shooting

Target shooting is the sport of firing at targets as an exercise in marksmanship. Rather than the ability to run a mile in 5 minutes or being able to lift your own body weight with one hand, the only requirements in target shooting are reasonable eyesight (glasses are fine, some of our best wear them), patience, concentration and tenacity. Gender is unimportant - all levels of the sport enjoy an even mix of both sexes, although women do have a distinct advantage in Air Rifle, and usually have a mild advantage in general. Antagonistic competitive attitudes are almost unheard-of in target shooting. The only person with whom you are really competing is yourself. This makes for a happy, relaxed atmosphere, even at the highest levels of the sport. A handicap classing system allows large competitions with competitors ranging from novices to old hands, which makes for great craic. Not that that means that it’s not a serious sport - target shooting has seventeen events in the olympics! Irish shooters have taken part in the last three olympics, and DURC has had at least one club member on the Irish National Squad for the past twenty years and we’ve had many club members compete internationally. This is an Air rifle tar-

Air Rifle Target

get. At a distance of ten metres, while standing and holding a rifle that weighs 5 kilograms and is over a meter long, you have to hit the white dot in the center sixty times in a row for a perfect score! Smallbore targets are larger, but because of the greater range over which they are shot, they appear to be the same relative size to the shooter. This is a sport that requires great patience and dedication to excel in, but which can at the same time provide a great deal of enjoyment to the casual shooter.

Equipment

The club provides all the equipment you’ll need. D.U.R.C. has equipment which is accurate enough for all levels from club competition right through to international competitions. Shooters wear a stiff double canvas jacket with non-slip pads and a glove. For Smallbore rifle, you also use a sling for the rifle, as it can weigh up to seven kilograms.

Shooting Jacket with sling and Shooting Glove

Competitions

D.U.R.C. has several internal com-

Air Rifle Open

petitions throughout the year, along with both hosting and attending national-level competitions. There are the Air and Smallbore Open competitions which are recognised by the N.T.S.A. for selection to the National Squad. We regularly compete against the UCD Rifle Club, in fact we have one competition, the Colours, where we compete against them alone. This is usually a popular so-

TCD & UCD Colours Teams

cial event as well as a serious rifle competition. We also enter the nationals each year in all disciplines, and for the last few years have usually done very well, often winning prizes. These competitions

are a great way to D.U.R.C. Squad at the UCD Open

meet the rest of the Irish shooting community, to gain experience and gather tips on shooting technique and so forth, not to mention the team spirit that goes with being a part of the D.U.R.C. squad.

About the club

DURC has been around for the last forty years. We’ve had members compete at all levels from internal club competitions to international world cups. We usually have between 300 and 500 members each year, with an equal mix of genders. We have an active calendar of events every year, both social and sporting. We shoot in several events : Olympic Air and Smallbore Rifle and Mexican Silouette rifle shooting. The DURC rifle

Firing lanes

range has facilities for shooting Smallbore rifles over 25 yards and Air rifles over 10 metres, the olympic standard for Air rifle. We also train in the Fassaroe Shooting Club’s 50 metre range, which is the Olympic standard distance for Smallbore rifle.

Seperate to the firing lanes, we

Clubhouse

also have two adjoining rooms which are used to prepare the rifles and equipment for shooting and for social activities, such as the annual Christmas Novelty match. Coffee, Tea and biscuit facilities are provided - but you have to wash your own mugs !

Learning to shoot safely

D.U.R.C. Range Officers are trained to coach new shooters to a very high standard of accuracy. Their main duty, however, is to ensure the safety of all those on the rifle range. On your first shooting detail, your Range Officer will explain to you the basic safety rules and how to shoot safely. Since our founding in 1962, D.U.R.C. has not had one single accident or injury (apart from the odd tall shooter banging his or her head off the roof of the range!). This is due to our committment to enjoying our sport in a safe manner, and reflects the Irish shooting community in general, where there hasn’t been an accident or injury in living memory.

How to find us

\begin{matrix}c \\ ◯ \end{matrix} Trinity College Dublin

The rifle range is located in Park Lane, behind the Civil Engineering building, as marked on the map. It’s the green door with DURC written on it, and the doorbell is to the right-hand side.